Sketch the graphs of y = x and y = sin(x). Use the bisection method to find an approximation within 10^{-5} to the first positive value of x with x = sin(2x).
Answer #1

y= x and Y = Sin(ex) y2x Now to gt- nt/en y y f(x)= x -in (2x) atb Bisecn mhod 2 iterakaon (o5)- -034 < ; f(j = o.0 o Yoot LifC0.75)- - o. < u; Jli D.1510.937 2 tertm - to.13 75) o1315, Yoot l ০.৭31 +। D.1697 2. iterda = o -03 >o /(0.9657) o.315 -6 -toratum -0.003< D fCo.749219) o75312, O 949219 f (o-45312) Youl w D 945312 0.941219 o47260 2. Th iteratm C0-47266) 0-60070 fC